# Quantum tomography with $\tau$ leptons at the FCC-ee Entanglement, Bell inequality violation, $\sin\theta_{W}$ and anomalous copulings

### v1

@@ 4 Outlook
T he FCC-ee can, after 4 years of operation at the energy of the $Z$-boson mass, test quantum entanglement and Bell inequality violation in $\tau$-lepton pairs with extraordinary precision. Already with the reduced sample we have used in our analysis the significance is well in excess of five standard deviations and it is expected to reach many times that value with the full expected dataset of $10^{9}$ events—provided the systematic uncertainties are under control. We believe that the results we have presented make a convincing case for including the quantum tomography of the $\tau$ leptons in the physics goals of the FCC-ee. The work also proves that the same quantum tomography technique can provide new and stringent bounds on the form factors entering the electroweak neutral current of the $\tau$ leptons.

### v2

@@ 4 Outlook
T he FCC-ee can, after 4 years of operation at the energy of the $Z$-boson mass, test quantum entanglement and Bell inequality violation in $\tau$-lepton pairs with extraordinary precision. Already with the reduced sample we have used in our analysis the significance is well in excess of five standard deviations. The same large significance is obtained with the full dataset of $10^{9}$ events—at which point the systematic uncertainty will dominate and any further progress will depend on better understanding this uncertainty. The same holds for the determination by quantum tomography of the Weinberg angle from the $\tau$-lepton polarizations and constraining the form factors entering the electroweak neutral current of the $\tau$ leptons. We believe that the results we have presented make a convincing case for including the quantum tomography of the $\tau$ leptons in the physics goals of the FCC-ee.